# Breaking Boundaries: Investigating the Effects of Model Editing on Cross-linguistic Performance

### v2

@@ 5.3 Evaluation metric
We evaluate the generated output from the edited models on four properties using two different metrics as follows.
Exact match: In this evaluation method, we systematically determine the accuracy of outputs generated through our model by checking for the presence of the ground truth within these outputs. The ground truth represents the exact, correct response expected from the model for a given input. An output is considered accurate (a ‘correct generation’) if it includes the ground truth, indicating the model’s ability to replicate the desired response accurately. Conversely, if the ground truth is not present in the output, it is classified as inaccurate (‘incorrect generation’). This binary evaluation approach provides a clear and direct measure of the model’s performance, focusing on its ability to produce outputs that are not only relevant but also correct as per the predefined standards.
Partial match: In case of partial match, the Levenshtein ratio is utilized as an alternative measure of textual similarity. This ratio is calculated by dividing the Levenshtein distance Levenshtein (1965), by the maximum length of either the ground truth text or the generated text. A critical threshold is set at an 80% Levenshtein ratio; outputs that do not contain the ground truth as a substring but surpass this ratio are deemed accurate. This approach allows for a nuanced assessment of the generated text’s correctness, accommodating minor deviations or errors that do not significantly detract from the overall meaning or accuracy of the generated output.
@@ 6 Results
ZsRE Models TowerInstruct Mistral OpenHathi Tamil-Llama Kan-Llama
Languages Properties ROME MEMIT ROME MEMIT ROME MEMIT ROME MEMIT ROME MEMIT
De Rel 0.48/0.59 0.25/0.30 0.51/0.62 0.38/0.47 - - - - - -
Gen 0.33/0.39 0.11/0.12 0.35/0.45 0.18/0.24 - - - - - -
Loc 0.00/0.01 0.00/0.01 0.01/0.02 0.01/0.03 - - - - - -
Port 0.02/0.02 0.00/0.00 0.08/0.10 0.02/0.04 - - - - - -
Es Rel 0.44/0.59 0.24/0.34 0.49/0.61 0.37/0.49 - - - - - -
Gen 0.30/0.40 0.16/0.20 0.35/0.45 0.22/0.29 - - - - - -
Loc 0.00/0.01 0.01/0.02 0.01/0.01 0.02/0.02 - - - - - -
Port 0.02/0.02 0.01/0.02 0.03/0.07 0.03/0.04 - - - - - -
It Rel 0.54/0.62 0.25/0.29 0.58/0.65 0.42/0.50 - - - - - -
Gen 0.35/0.43 0.16/0.20 0.42/0.48 0.25/0.31 - - - - - -
Loc 0.00/0.00 0.00/0.01 0.00/0.02 0.01/0.02 - - - - - -
Port 0.01/0.02 0.02/0.03 0.07/0.08 0.01/0.03 - - - - - -
Fr Rel 0.51/0.59 0.27/0.35 0.52/0.63 0.40/0.50 - - - - - -
Gen 0.28/0.35 0.14/0.17 0.40/0.50 0.19/0.27 - - - - - -
Loc 0.00/0.01 0.00/0.02 0.01/0.02 0.01/0.02 - - - - - -
Port 0.03/0.05 0.03/0.03 0.06/0.09 0.04/0.06 - - - - - -
Hi Rel - - - - 0.03/0.06 0.20/0.33 - - - -
Gen - - - - 0.01/0.04 0.19/0.28 - - - -
Loc - - - - 0.01/0.01 0.00/0.01 - - - -
Port - - - - 0.00/0.00 0.03/0.03 - - - -
Ta Rel - - - - - - 0.06/0.08 0.16/0.21 - -
Gen - - - - - - 0.03/0.04 0.10/0.14 - -
Loc - - - - - - 0.00/0.00 $0$
Port - - - - - - 0.00/0.00 0.01/0.01 - -
Kn Rel - - - - - - - - 0.16/0.21 0.05/0.07
Gen - - - - - - - - 0.08/0.17 0.05/0.05
Loc - - - - - - - - 0.00/0.00 0.00/0.00
Port - - - - - - - - 0.00/0.01 0.00/0.00
Table 2: Comparison of reliability, generalization, locality, and portability scores across various language models, evaluated for the ZsRE dataset under same language edit and inference settings. Each cell shows results in the format: exact match/partial match. The highest score presented in bold whereas second highest presented in underlined.
CounterFact Models TowerInstruct Mistral OpenHathi Tamil-Llama Kan-Llama
Languages Properties ROME MEMIT ROME MEMIT ROME MEMIT ROME MEMIT ROME MEMIT
De Rel 0.48/0.53 0.40/0.46 0.50/0.56 0.54/0.61 - - - - - -
Gen 0.25/0.27 0.13/0.17 0.23/0.27 0.22/0.23 - - - - - -
Loc 0.20/0.21 0.19/0.22 0.23/0.25 0.26/0.28 - - - - - -
Port 0.00/0.00 0.00/0.00 0.03/0.03 0.03/0.04 - - - - - -
Es Rel 0.51/0.56 0.40/0.48 0.57/0.62 0.56/0.60 - - - - - -
Gen 0.26/0.29 0.18/0.22 0.25/0.29 0.21/0.26 - - - - - -
Loc 0.22/0.24 0.17/0.17 0.24/0.27 0.25/0.27 - - - - - -
Port 0.00/0.00 0.00/0.00 0.03/0.03 0.03/0.04 - - - - - -
It Rel 0.45/0.50 0.35/0.40 0.47/0.58 0.44/0.49 - - - - - -
Gen 0.23/0.27 0.19/0.20 0.25/0.35 0.21/0.23 - - - - - -
Loc 0.20/0.21 0.20/0.20 0.24/0.36 0.28/0.29 - - - - - -
Port 0.01/0.02 0.01/0.02 0.03/0.11 0.04/0.04 - - - - - -
Fr Rel 0.50/0.53 0.45/0.49 0.49/0.55 0.51/0.59 - - - - - -
Gen 0.28/0.31 0.19/0.22 0.28/0.31 0.26/0.27 - - - - - -
Loc 0.23/0.23 0.19/0.21 0.20/0.36 0.25/0.26 - - - - - -
Port 0.01/0.01 0.01/0.01 0.01/0.12 0.03/0.04 - - - - - -
Hi Rel - - - - 0.56/0.66 0.02/0.03 - - - -
Gen - - - - 0.27/0.34 0.03/0.03 - - - -
Loc - - - - 0.26/0.31 0.03/0.03 - - - -
Port - - - - 0.02/0.02 0.00/0.01 - - - -
Ta Rel - - - - - - 0.00/0.00 0.01/0.01 - -
Gen - - - - - - 0.00/0.00 $0$
Loc - - - - - - 0.01/0.01 0.01/0.02 - -
Port - - - - - - 0.00/0.00 $0$
Kn Rel - - - - - - - - 0.01/0.01 0.00/0.00
Gen - - - - - - - - 0.00/0.01 0.00/0.00
Loc - - - - - - - - 0.02/0.02 0.03/0.03
Port - - - - - - - - 0.00/0.00 0.00/0.00
Table 3: Comparison of reliability, generalization, locality, and portability scores across various language models, evaluated for the CounterFact dataset under English edit and target language inference settings. Each cell shows results in the format: exact match/partial match.
ZsRE Models TowerInstruct Mistral OpenHathi Tamil-Llama Kan-Llama
Languages Properties ROME MEMIT ROME MEMIT ROME MEMIT ROME MEMIT ROME MEMIT
De Rel 0.24/0.28 0.10/0.14 0.34/0.45 0.14/0.18 - - - - - -
Gen 0.18/0.23 0.12/0.14 0.26/0.35 0.14/0.16 - - - - - -
Loc 0.00/0.01 0.00/0.02 0.01/0.02 0.01/0.03 - - - - - -
Port 0.02/0.02 0.02/0.02 0.06/0.07 0.02/0.03 - - - - - -
Es Rel 0.24/0.29 0.12/0.14 0.39/0.48 0.19/0.26 - - - - - -
Gen 0.18/0.25 0.09/0.11 0.33/0.41 0.14/0.21 - - - - - -
Loc 0.00/0.01 0.01/0.02 0.01/0.02 0.02/0.02 - - - - - -
Port 0.02/0.03 0.01/0.01 0.04/0.06 0.04/0.05 - - - - - -
It Rel 0.24/0.29 0.12/0.14 0.31/0.34 0.23/0.27 - - - - - -
Gen 0.17/0.22 0.11/0.13 0.26/0.32 0.18/0.21 - - - - - -
Loc 0.00/0.00 0.00/0.01 0.00/0.02 0.01/0.02 - - - - - -
Port 0.01/0.02 0.02/0.02 0.07/0.08 0.01/0.01 - - - - - -
Fr Rel 0.22/0.26 0.12/0.17 0.36/0.44 0.23/0.28 - - - - - -
Gen 0.15/0.21 0.08/0.10 0.29/0.33 0.16/0.21 - - - - - -
Loc 0.00/0.01 0.00/0.02 0.01/0.03 0.01/0.02 - - - - - -
Port 0.02/0.02 0.02/0.02 0.06/0.09 0.04/0.05 - - - - - -
Hi Rel - - - - 0.03/0.03 0.03/0.06 - - - -
Gen - - - - 0.03/0.03 0.04/0.08 - - - -
Loc - - - - 0.00/0.00 0.00/0.01 - - - -
Port - - - - 0.00/0.00 0.01/0.01 - - - -
Ta Rel - - - - - - 0.00/0.00 0.01/0.01 - -
gen - - - - - - 0.00/0.00 $0$
Loc - - - - - - 0.00/0.00 $0$
Port - - - - - - 0.00/0.00 $0$
Kn Rel - - - - - - - - 0.00/0.01 0.00/0.01
gen - - - - - - - - 0.00/0.00 0.00/0.00
Loc - - - - - - - - 0.00/0.00 0.00/0.00
Port - - - - - - - - 0.00/0.00 0.00/0.00
Table 4: Comparison of reliability, generalization, locality, and portability scores across various language models, evaluated for the ZsRE dataset under English edit and target language inference settings. Each cell shows results in the format: exact match/partial match.
@@
Table 5: Comparison of reliability, generalization, locality, and portability scores across the merge model for the three Indic languages, evaluated using the CounterFact and ZsRE dataset under each language for itself and others. Each cell shows results in the format: exact match/partial match. The highest score presented in bold whereas second highest presented in underlined. The highest score presented in bold whereas second highest presented in underlined.
@@ 6.1 Self edit - self inference perspective
CounterFact dataset: In our evaluations of the model performance for the CounterFact dataset, we observe marked variations across different languages and metrics in Table 1, illustrating significant challenges in multilingual adaptability and contextual understanding. For instance, German language tests show that models like TowerInstruct and Mistral achieve good reliability scores (ROME at 0.83 and MEMIT at 0.73 for TowerInstruct; the same scores are at 0.83 and 0.73 respectively for Mistral). These scores illustrate good model performance in understanding the contextual nuances of German. However, generalization and locality score are less impressive (TowerInstruct at 0.27 and 0.22 on ROME for generalization and locality respectively), indicating difficulties in applying the learned information across broader contexts and different locales within the German language. Similar patterns are observed in Spanish and Italian. In Spanish, TowerInstruct reaches a reliability score of 0.82 for ROME and 0.70 for MEMIT; for Mistral the reliability scores are 0.81 for ROME and 0.78 for MEMIT, suggesting decent grasp of Spanish contexts. However, the generalization scores remain below 0.35 for ROME and locality scores do not exceed 0.29 for MEMIT for any model. Despite TowerInstruct showing a relatively high reliability in Italian with a ROME at 0.87 and MEMIT at 0.74, the generalization and locality scores remain low (highest being 0.35 on ROME and 0.28 on MEMIT for Mistral). In case of the three Indic languages the discrepancies become even more pronounced. OpenHathi, for example, shows a drastic drop in Hindi, with a ROME reliability of just 0.02 and a MEMIT of 0.45, indicating almost no comprehension of the language nuances. Tamil-Llama and Kan-Llama also display low scores across all properties. The highest reliability achieved is 0.21 for ROME for Kan-Llama and 0.48 for MEMIT in case of Tamil-Llama, which highlights the limitations in these language models. Portability scores are consistently low across all languages, models, and metrics, demonstrating a significant gap in model training as it fails to effectively account for diverse linguistic structures and cultural contexts.
Id Category Examples Possible solution
1 Lexical ambiguity
@@ 7 Idioms
8 Phonetic/orthographic errors
@@ 9 Morphological errors
10 Pragmatic errors
French: Using ‘tu’ (informal you) instead of ‘vous’ (formal or plural you) in a formal context can be seen as rude or too casual.
Understanding cultural norms
Table 6: A systematic categorization of errors encountered in the process of multilingual knowledge editing, spanning lexical, syntactic, semantic, cultural, and contextual ambiguities, among others. Each category is illustrated with examples from English, French, German, Italian, and Spanish, showcasing the multifaceted challenges of maintaining consistency and accuracy in knowledge representation across languages.
ZsRE dataset: In case of ZsRE dataset (see Table 2) German shows moderate performance in reliability with scores like 0.48 on ROME and 0.25 on MEMIT for TowerInstruct. The generalization (0.33 for ROME) and locality scores ($\sim 0$) are also very poor. These results indicate substantial deficiencies in capturing language-specific details and generalizing learned information. Spanish fares slightly better in reliability, achieving up to 0.49 on ROME with TowerInstruct and Mistral, but like German, faces challenges in generalization and locality, with the best generality reaching only 0.35 and locality remaining near zero. Italian (It) generally scores higher in reliability, particularly with Mistral reaching 0.58 on ROME, though it too struggles with generality and locality. French exhibits a similar trend, with reliability scores reaching up to 0.52 for ROME with Mistral and both generalization and locality scores remaining low. Performance markedly drops for the three Indic languages. For instance, Hindi’s highest reliability is just 0.03 for ROME, while Tamil and Kannada only achieve maximum reliability scores of 0.06 and 0.16 respectively for ROME. Across all languages, portability scores are low, reflecting limited adaptability and the challenge of transferring learned capabilities from one linguistic context to another.
@@ 6.2 English edit - self inference perspective
CounterFact dataset: In German, the reliability scores for models such as TowerInstruct and Mistral suggest moderate effectiveness, with ROME around 0.48 and MEMIT around 0.40 (see Table 3). However, their generalization and locality scores reveal limitations in the models’ ability to generalize and localize content effectively with scores not exceeding 0.25 and 0.26 respectively. For Spanish, there is a noticeable improvement in reliability, with ROME scores for Mistral reaching 0.57, and a slight improvement in generalization and locality metrics compared to German. Italian and French show similar trends, with reliability scores peaking at 0.47 for Mistral in Italian and 0.49 in French; the generalization and locality scores are still lower. For Tamil and Kannada the reliability are exceptionally low. In fact, in case of Tamil this score is 0 for ROME and 0.01 for MEMIT. Comparatively for Hindi the reliability scores are quite good with 0.56 for ROME. However the portability and generalization scores are again very poor.
ZsRE dataset: For languages such as German and Spanish, the models display moderate reliability with Mistral, achieving ROME scores up to 0.34 and 0.39 respectively, and MEMIT scores of 0.14 and 0.19 respectively (see Table 4). However, the scores significantly drop for locality and portability, showing that while the models can identify relevant relationships, they struggle to generalize and adapt to the specific linguistic nuances of these languages. The trends are similar in Italian and French, where reliability scores are moderate while locality and generalization scores are poor. Further, for the Indic languages, the score are exceedingly low for all the properties indicating the stark gap in performance highly resource scarce languages. .
@@ 6.3 Merged model perspective
The Table 5 presents performance metrics for the merged model, showcasing how it handles edits across various languages, with columns representing the inferencing language and rows indicating the respective editing language. When both editing and inferencing are done in English the reliability scores for the CounterFact dataset are quite high with ROME and MEMIT respectively reaching 0.73 and 0.95. However, the performance sharply declines when editing in English and inferencing is in Hindi, Tamil, and Kannada with scores nearly zero, highlighting a stark limitation in the model’s cross-lingual capabilities. This trend is consistently observed across both datasets. Editing in languages like Hindi, Tamil, or Kannada results in uniformly poor outcomes across all four properties regardless of the inferencing language, which indicates profound inadequacies in the model’s ability to generalize and adapt across linguistic barriers. This pattern emphasizes the critical need for advancing multilingual model adaptability. The current findings suggest that while the model operates effectively within the confines of the same linguistic environment, its performance deteriorates dramatically across linguistic boundaries, especially from lesser-resourced languages. Such insights advocate for a significant enhancement in training approaches, aiming to foster robust multilingual support and ensure that models are truly multilingual in functionality, proficiently managing edits and inferencing across a diverse linguistic landscape.
@@ 7 Error analysis
In Table 6 we show the different types of linguistic errors encountered during the translation and editing process. The errors are categorised based on the different types of ambiguities and sheds light on how future models should strengthened by carefully harnessing techniques to tackle these errors.
@@ 8 Conclusion
Appendix A Hyperparameters

### v3

@@ 5.3 Evaluation metric
We evaluate the edited models using two metrics:
Exact match: Here accuracy is determined by checking if the ground truth is present in the model’s output. Outputs containing the exact expected response are classified as correct, while others are deemed incorrect, providing a binary measure of performance.
Partial match: The Levenshtein ratio Levenshtein (1965) measures textual similarity, calculated as the Levenshtein distance divided by the maximum text length. Outputs surpassing an 80% ratio but not containing the ground truth as a substring are considered accurate, allowing for minor acceptable deviations.
@@ 6 Results
Table 4: Comparison of scores across the merged model for three Indic languages, evaluated using the CounterFact and ZsRE datasets for each language and others. Highest scores are in bold, and second-highest are underlined. Values represent Exact Match/Partial Match results.
@@ 6.1 Self edit - self inference perspective
CounterFact dataset: In our evaluations of the model performance for the CounterFact dataset, we observe marked variations across different languages and metrics in Table 1, illustrating significant challenges in multilingual adaptability and contextual understanding. For instance, German language tests show that models like TowerInstruct and Mistral achieve good reliability scores (ROME at 0.83 and MEMIT at 0.73 for TowerInstruct; the same scores are at 0.83 and 0.73 respectively for Mistral). These scores illustrate good model performance in understanding the contextual nuances of German. However, generalization and locality score are less impressive (TowerInstruct at 0.27 and 0.22 on ROME for generalization and locality respectively), indicating difficulties in applying the learned information across broader contexts and different locales within the German language. Similar patterns are observed in Spanish and Italian. In Spanish, TowerInstruct reaches a reliability score of 0.82 for ROME and 0.70 for MEMIT; for Mistral the reliability scores are 0.81 for ROME and 0.78 for MEMIT, suggesting decent grasp of Spanish contexts. However, the generalization scores remain below 0.35 for ROME and locality scores do not exceed 0.29 for MEMIT for any model. Despite TowerInstruct showing a relatively high reliability in Italian with a ROME at 0.87 and MEMIT at 0.74, the generalization and locality scores remain low (highest being 0.35 on ROME and 0.28 on MEMIT for Mistral). In case of the three Indic languages the discrepancies become even more pronounced (See Table 3). OpenHathi, for example, shows a drastic drop in Hindi, with a ROME reliability of just 0.02 and a MEMIT of 0.45, indicating almost no comprehension of the language nuances. Tamil-Llama and Kan-Llama also display low scores across all properties. The highest reliability achieved is 0.21 for ROME for Kan-Llama and 0.48 for MEMIT in case of Tamil-Llama, which highlights the limitations in these language models. Portability scores are consistently low across all languages, models, and metrics, demonstrating a significant gap in model training as it fails to effectively account for diverse linguistic structures and cultural contexts.
Category Examples Possible solution
Lexical ambiguity
@@
Phonetic/orthographic errors
@@
Pragmatic errors French: Using ‘tu’ (informal you) instead of ‘vous’ (formal or plural you) in a formal context can be seen as rude or too casual. Understanding cultural norms
Table 5: Categorization of multilingual knowledge editing errors, including lexical, syntactic, semantic, cultural, and contextual ambiguities, with examples from English, French, German, Italian, and Spanish, highlighting challenges in cross-lingual consistency and accuracy.
ZsRE dataset: In case of ZsRE dataset (see Table 1) German shows moderate performance in reliability with scores like 0.48 on ROME and 0.25 on MEMIT for TowerInstruct. The generalization (0.33 for ROME) and locality scores ($\sim 0$) are also very poor. These results indicate substantial deficiencies in capturing language-specific details and generalizing learned information. Spanish fares slightly better in reliability, achieving up to 0.49 on ROME with TowerInstruct and Mistral, but like German, faces challenges in generalization and locality, with the best generality reaching only 0.35 and locality remaining near zero. Italian (It) generally scores higher in reliability, particularly with Mistral reaching 0.58 on ROME, though it too struggles with generality and locality. French exhibits a similar trend, with reliability scores reaching up to 0.52 for ROME with Mistral and both generalization and locality scores remaining low. Performance markedly drops for the three Indic languages (See Table 3). For instance, Hindi’s highest reliability is just 0.03 for ROME, while Tamil and Kannada only achieve maximum reliability scores of 0.06 and 0.16 respectively for ROME. Across all languages, portability scores are low, reflecting limited adaptability and the challenge of transferring learned capabilities from one linguistic context to another.
@@ 6.2 English edit - self inference perspective
CounterFact dataset: In German, the reliability scores for models such as TowerInstruct and Mistral suggest moderate effectiveness, with ROME around 0.48 and MEMIT around 0.40 (see Table 2). However, their generalization and locality scores reveal limitations in the models’ ability to generalize and localize content effectively with scores not exceeding 0.25 and 0.26 respectively. For Spanish, there is a noticeable improvement in reliability, with ROME scores for Mistral reaching 0.57, and a slight improvement in generalization and locality metrics compared to German. Italian and French show similar trends, with reliability scores peaking at 0.47 for Mistral in Italian and 0.49 in French; the generalization and locality scores are still lower. For Tamil and Kannada the reliability are exceptionally low (See Table 3). In fact, in case of Tamil this score is 0 for ROME and 0.01 for MEMIT. Comparatively for Hindi the reliability scores are quite good with 0.56 for ROME. However the portability and generalization scores are again very poor.
@@
[…]
@@ 6.3 Merged model perspective
Table 4 presents performance metrics for the merged model, with columns representing inferencing languages and rows indicating editing languages. Editing and inferencing in English yield high reliability scores on the CounterFact dataset (ROME: 0.73, MEMIT: 0.95). However, performance drops to near zero when editing in English and inferencing in Hindi, Tamil, or Kannada, exposing the model’s cross-lingual limitations. Editing in Hindi, Tamil, or Kannada consistently results in poor outcomes across all properties, regardless of the inferencing language. This highlights the model’s inability to generalize across linguistic barriers and underscores the need for improved multilingual adaptability. The findings reveal that while the model performs well within the same linguistic environment, its performance deteriorates significantly across lesser-resourced languages, necessitating enhanced training approaches for robust multilingual support.
@@ 7 Error analysis
In Table 5 we show the different types of linguistic errors encountered during the translation and editing process. The errors are categorised based on the different types of ambiguities and sheds light on how future models should strengthened by carefully harnessing techniques to tackle these errors. More details are available in Appendix B.
@@ 8 Conclusion
9 Limitations
Despite the promising results, our study has several limitations. The variability in performance across different languages highlights the inherent challenges in achieving true multilingual consistency, with models exhibiting substantial difficulties in generalizing and localizing edits, particularly in low-resourced languages such as Hindi, Tamil, and Kannada. This discrepancy indicates a need for more inclusive and representative training datasets that encompass a wider range of linguistic and cultural contexts. Additionally, our focus on decoder-only models limits the generalizability of our findings to other types of language models, such as encoder-decoder architectures. The relatively low portability scores across all languages further indicate that current models struggle to transfer learned knowledge effectively from one linguistic context to another, especially in cross-lingual edits where modifications in one language often fail to translate accurately into another. Moreover, the merging of models, while showing some promise, does not consistently improve reliability, locality, or generalization metrics, suggesting that further research is needed to optimize these approaches.
10 Ethical consideration
Our research raises ethical concerns regarding linguistic equity and cultural sensitivity. Disparities in model performance could reinforce existing linguistic inequities, limiting access to AI technologies for speakers of low-resourced languages. Future model development must include diverse languages and dialects to promote equity. Additionally, errors related to cultural ambiguity and idiomatic expressions can lead to misinterpretations or offensive content, necessitating robust evaluation frameworks to ensure cultural sensitivity. Privacy and security risks are also significant, as models may inadvertently reveal sensitive information during knowledge editing processes. Researchers must prioritize user privacy and implement stringent data protection measures to prevent misuse of personal data, ensuring AI technologies are effective and equitable for all users.
11 Potential risk
LLMs can be used for harmful content generation and misinformation spread. The prompts used and generated in this work can be misused to generate harmful content.
Appendix A Model selection
Mistral-7B-Instruct-v0.2101010https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2: The model was developed by Jiang et al. (2023) and supports multilinguality [^11] https://encord.com/blog/mistral-large-explained/. It is designed around the causal language modeling framework. We shall refer to this model as Mistral.
TowerInstruct-7B-v0.2121212https://huggingface.co/Unbabel/TowerInstruct-7B-v0.2: This model Alves et al. (2024) has been developed on top of LLaMA2 Touvron et al. (2023) architecture and supports multilinguality including English, German, French, Spanish, Chinese, Portuguese, Italian, Russian, Korean, and Dutch. We shall refer to this model as TowerInstruct.
Tamil-llama-7b-base-v0.1141414https://huggingface.co/abhinand/tamil-llama-7b-base-v0.1: This is a sophisticated model Balachandran (2023) developed specifically for bilingual tasks in Tamil and English, leveraging a 7 billion parameter causal language modeling framework. We shall refer to this model as Tamil-Llama.
Appendix B Error analysis
Lexical ambiguity Lexical ambiguity occurs when a word has multiple meanings, leading to confusion without context. For instance, the English word "crane" can refer to a bird or construction equipment, a distinction crucial for accurate knowledge representation.
Syntactic ambiguity Syntactic ambiguity arises from sentence structures that can be interpreted in multiple ways. An example is the English sentence "Visiting relatives can be boring," which could imply either the act of visiting relatives is boring or that the relatives being visited are boring. Resolving these ambiguities requires advanced parsing techniques and an understanding of the specific language’s syntax to ensure accurate interpretation.
Semantic Ambiguity Errors Semantic ambiguity pertains to the uncertainty of meaning within a sentence or phrase. For example, "He gave her a ring" could mean a telephone call or presenting a piece of jewelry. Multilingual systems need to discern the intended meaning based on semantic cues and the broader context, a challenging task given the subtlety of cues and cultural specificities in language use.
Cultural and Contextual Errors These errors occur when language processing fails to account for cultural idioms or context-specific meanings. Phrases like "Piece of cake" in English, meaning something easy, can be misunderstood if taken literally or translated directly into another language without considering idiomatic expressions. Handling these requires deep cultural knowledge and contextual understanding beyond linguistic analysis.
Translation Errors Translation errors emerge when converting text from one language to another, often leading to loss of meaning or inaccuracies. These can be particularly problematic in knowledge editing, where precision is paramount. For example, translating idiomatic expressions or culturally specific terms often requires not just a direct translation but a reinterpretation in the target language.
Named Entity Recognition (NER) Errors NER errors involve the incorrect identification or classification of proper nouns in text. For instance, distinguishing between "Rio" as a river or the city of Rio de Janeiro in Spanish requires contextual analysis. Accurate NER is essential for knowledge databases to correctly link information to entities, demanding sophisticated language models that can navigate these nuances.
Idiomatic Expression Errors Errors in understanding or translating idiomatic expressions can significantly alter the intended meaning. For example, the Italian idiom "Tra il dire e il fare c’è di mezzo il mare" illustrates the difference between saying and doing, a concept that might be lost if translated literally. Addressing these requires an in-depth understanding of both the source and target languages’ idioms.
Phonetic and Orthographic Errors These errors occur with words that sound similar (homophones) or are spelled similarly (homographs) but have different meanings. For instance, "their," "there," and "they’re" in English. Multilingual systems must accurately identify and apply the correct form based on context, a challenging task that often requires human-like understanding of language.
Morphological Errors Morphological errors refer to the misuse of word forms, affecting the grammatical structure and potentially changing the meaning of sentences. German’s gender-specific articles—der, die, das—offer a prime example, where incorrect usage can confuse readers and misrepresent information. Overcoming these demands a robust grasp of linguistic rules and the flexibility to apply them in diverse contexts.
Pragmatic Errors Pragmatic errors involve the misuse or misunderstanding of language in social context, such as politeness or formality levels. An example is the inappropriate use of "tu" (informal) and "vous" (formal or plural) in French, which can significantly affect the tone and perceived respectfulness of an interaction. Addressing these requires sensitivity to cultural norms and the social dynamics of language, highlighting the complexity of human communication and the challenges in replicating these nuances in AI systems.
Appendix C Hyperparameters